OpenStudy (anonymous):

Find the X interceps using the factor method, write the equation of the aqis of symmetry.vertex,state if its max or min............O.o I DONT UNDERSTAND

OpenStudy (anonymous):

-x^2+3x+10

OpenStudy (anonymous):

-x^2+5x-2x+10=(5-x)(x+2)

OpenStudy (anonymous):

therefore x-intercepts are 5 and -2

OpenStudy (anonymous):

In step 2 where did u get 5-x and x+2 from?

OpenStudy (anonymous):

-x^2+5x-2x+10=x(5-x)+2(5-x)=(5-x)(x+2)

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

not even a little bit......

OpenStudy (anonymous):

i factorized -x^2+3x+10=(5-x)(x+2). Suppose ur given function is y. then at x-intercepts y=0...therefore (5-x)(x+2)=0..i.e. x=5,-2 :)

OpenStudy (anonymous):

how do you factorize?????????

OpenStudy (anonymous):

ok...split 3x as 5x-2x then -x^2+5x-2x+10

OpenStudy (anonymous):

or u can think it of as: put x=-2 and x=5 in ur equation ..u get zero for these two values..so x-intercepts are -2 and 5..

OpenStudy (anonymous):

I think you should learn factorizing before doing coordinate geometry :)

OpenStudy (anonymous):

ok i got it thank you

OpenStudy (anonymous):

now how do u find vertex?

OpenStudy (anonymous):

r u sure u want to proceed..cause u need great deal of factoring!!

OpenStudy (anonymous):

take y=-x^2+3x+10 then u get by arranging and factoring (x-3/2)^2=4(-1/4)(y-49/4)

OpenStudy (anonymous):

Phil7780 go to my youtbe link on factorising quadratics nd let me know if it helps. http://www.youtube.com/watch?v=PYTo3b5JO7Y